# Semialgebraic Neural Networks: From roots to representations

### v1

@@ 1 Introduction
Many classical numerical algorithms compute semialgebraic functions—functions whose graphs are defined by finitely many polynomial equalities and inequalities. Such functions include anything computable using finitely many real variables and finitely many operations addition, subtraction, multiplication, division, extraction of roots, and branching if statements with polynomial decision boundaries. Familiar primitives from numerical linear algebra, such as solving linear systems, Schur complements, and extraction of real eigenvalues are also semialgebraic, as are the solution operators of optimization problems with polynomial objective functions and constraints Press et al. (2007). Polynomial functions of $n$-th roots $n]{\cdot}$ are widely used in computation of eigenvalues of non-symmetric matrices and in the perturbation theory of eigenvalues of linear operators Kato (1995). Also, Finite Element approximations of partial differential equations lead to solving linear systems which solutions are semialgebraic functions Beltzer (1990). Due to their ubiquity, it is likely that the space of semialgebraic functions provides good inductive bias for operator learning with neural networks. This ansatz leads to our central question: Is it possible to design a neural network architecture capable of exactly computing any semialgebraic function?
In this paper, we combine techniques from classical numerical analysis, namely polynomial homotopy continuation methods for root finding, with ReLU activation functions to create neural networks capable of exactly computing any bounded semialgebraic function. The starting point for our idea is the observation that, under mild conditions, the graph of a semialgebraic function $F$ given by the relation $F(x)=y$ can be encoded as the kernel of a continuous piecewise polynomial $G$, i.e., the set of $(x,y)$ where $G(x,y)=0$ (Section 2). Therefore, in principle, we can use a neural network combining polynomials and ReLU activations to learn $G$, then append a root-finding procedure such as Newton’s method to compute $F(x)=\texttt{root}(G(x,\cdot))=y$ in a manner similar to Bai et al. (2019). However, this straightforward approach has numerous practical difficulties: there is no guarantee such a neural network would have a root $y$ for every $x$, roots may not be unique, the root-finding procedure may never converge, etc. To remedy these issues, we take inspiration from homotopy continuation methods for root finding. We construct a function $H(x,y,s)$ that continuously deforms from a simple function $G_{0}$ to the target function $G$:
@@ 1.2 Our contribution
We introduce the SANN architecture, and provided an exact characterization of the functions it represents. Our key contributions are as follows:
@@
We also argue SANNs are naturally suited to a wide variety of difficult problems, with application to solving linear systems (Section 5), nonlinear inverse problems (Section F), general optimization problems (Section G.1), and transformers (Section G.3). We also demonstrate through numerical experiments that SANNs can be trained using standard techniques (Sections 5.2 and F.1), and discuss future research directions (Section 6).
@@ 2.1 Inf-sup definable piecewise polynomials
Our networks are constructed from piecewise polynomials, formed by combining polynomials with ReLU activations. Using these components, we can represent the pointwise minimum and maximum of any finite collection of polynomials. In mathematical literature, such structures are referred to as lattices. Instead of “min” and “max,” the terms “inf” and “sup” are traditionally used in lattice theory; since our constructions rely heavily on lattice theory, we adopt this terminology here. Notably, we do not use the word “lattice” to refer to a grid-like structure (e.g., $\mathbb{Z}^{d}$), but rather a set of functions that is closed under the “min” and “max” operations.
@@
def-isd Let $D\subseteq\mathbb{R}^{m}$. $ISD(D)$ is the lattice generated by the polynomials $\mathbb{R}[x_{1},\dots,x_{m}]$, that are viewed as functions $D\to\mathbb{R}$, and the $\min$ and $\max$ operations. The vectors of $n$ $ISD(D)$ functions are denoted $ISD(D,\mathbb{R}^{n})$. For any non-negative integer $k$, $ISD^{k}(D,\mathbb{R}^{n})$ are the $k$-times differentiable functions in $ISD(D,\mathbb{R}^{n}).$
The shorthand “$f$ is $ISD$” means $f\in ISD(D,\mathbb{R}^{n})$ for some appropriate $D$ and $n$. Our nomenclature is taken from Mahé (1984, 2007).
@@
It is currently unknown whether every continuous real piecewise polynomial is ISD; this is the famous “Pierce–Birkhoff conjecture”. ISD functions suffice for our purpose to build semialgebraic neural networks, regardless whether the Pierce–Birkhoff conjecture holds or not.
@@ 2.3 Semialgebraic functions as kernels of ISD functions
prop-semialgebraic-sets-as-kernels $S\subset\mathbb{R}^{m}$ is a closed semialgebraic set if and only if there exists $f\in ISD^{1}(\mathbb{R}^{m},\mathbb{R}_{\geq 0})$ such that $\ker(f)=S$, where $\ker(f)=\{x\in\mathbb{R}^{m}:f(x)=0\}$.
@@ 3.2 SANN architecture
$z_{j+1}=\hbox{ODE-step}\left(\dot{z}_{j},z_{j},\frac{j}{N}\right):=z_{j}+\frac{1}{N}\dot{z}_{j}$ (2)
@@
$\dot{z}_{j}=\hbox{clamp-sol}(M(x,z_{j},\frac{j}{N}),b(x,z_{j},\frac{j}{N})),$ (3)
$z_{0}=0,$ (4)
where matrix $M(x,z,s)$ and vector $b(x,z,s)$ are the output of an ISD network ${\mathcal{N}}(x,z,s)$. Moreover, the function clamp-sol is defined
@@
The function $\mathrm{clamp}$ operates component-wise on each element of a vector, and is defined
$\mathrm{clamp}(a,low,high):=\begin{cases}a,&a\in[low,high]\\ low,&a<low\\ high&a>high.\end{cases}$ (6)
These functions guarantee the iteration (3)–(4) is well-defined even when $\mathcal{N}$ produces singular or nearly-singular output matrices $M$. Lastly, $\Pi$ projects onto the first $n$ components of its input.
The iteration (3)–(4) is the Euler finite-difference scheme that approximates the solution of the ODE
$\dot{z}(s)=\hbox{clamp-sol}(M(x,z(s),s),b(x,z(s),s)),$ (7)
@@ 3.3 SANN architecture: Algorithmic presentation
In Algorithm 1, $\texttt{ODESolve}(\dot{z},z(0))$ outputs $z_{N}$ according to the finite difference scheme in formulas (2)–(4), or another numerical solver chosen by the programmer. This interface reflects popular libraries for numerically solving ODEs, e.g. Malengier et al. (2018); Kidger (2021).
@@
Our expressivity proofs below refer to the true solution $z(1)$ to the ODE (7)-(8) rather than the numerical solution $z_{N}.$ The numerical ODE solver may introduce approximation error into SANNs, even for semialgebraic functions, but this approximation error is of a fundamentally different character than that which normally arises in neural networks, and the mitigation strategies will be different for SANNs. For example, increasing the number of steps $N$ will decrease the approximation error without increasing the number of parameters in the network. On the other hand, increasing the number of parameters in $\mathcal{N}$ does not decrease the approximation error, as it generally would for other architectures. Furthermore, SANNs may be constructed using other numerical strategies besides the finite-difference scheme (2). Changing this scheme does not affect our expressivity proofs in this paper, but there may be other implications in practice.
@@ 3.4 Training
In the supervised learning regime, we are given training data $\{(x^{(\ell)},y^{(\ell)}):\ \ell=1,2,\dots,L\}$, $y^{(\ell)}=F(x^{(\ell)})$ where $F$ is the target semialgebraic function. The output of a SANN with parameter $x$ is $y_{N}(x):=\Pi z_{N}$ and the intermediate update directions given in equation 3 are $\dot{y}_{j}(x):=\dot{z}_{j}$. In our experiments, we found it helpful to include a “direction loss” term to train each $y_{j}$ for $j=1,\dots,N$:
@@ 4 Expressivity of SANNs
In this section, we show that SANNs are capable of exactly representing any bounded semialgebraic function. We start by showing that every continuous semialgebraic function can be evaluated by a particular homotopy continuation method that is computable by iteratively solving the ODE (7)-(8) with different parameter functions $\mathcal{N}$. We then extend this result to cover discontinuous semialgebraic functions by first separating each connected component of the graph, applying results for the continuous case to obtain homotopies that represent each piece, then glueing these homotopies together in a manner consistent with ODE (7)-(8) to obtain the SANN. Additional details, including background in semialgebraic geometry and full proofs of the theorems we present here, are in Appendices D and E.
@@ 4.1 Continuous semialgebraic functions
The next theorem shows the family of homotopies in equation (11) is capable of representing any continuous bounded semialgebraic function $F$ for almost every choice of $a$; more precisely every $a$ except possibly on a semialgebraic set of dimension less than $n$ c.f. Transversality theorem in D.1. The requirement that $F$ be bounded is always satisfied when the domain of continuous $F$ is compact, which is a common and reasonable assumption in the context of machine learning with finite training data. The proof is in appendix E.
@@
thm-continuous-semialgebraic-functions-computable-by-homotopy Let $F:\mathbb{R}^{m}\to U_{F}$ be a given continuous semialgebraic function. For every ${x\in\mathbb{R}^{m}}$ there exists $H_{x,a}$ with the form equation 11 such that for almost every ${a\in U_{F}}$, the following hold:
@@
Theorem LABEL:thm-continuous-semialgebraic-functions-computable-by-homotopy is used as follows: When $t=0$ the function $y\to H_{x,a}(y,0)$ has the root $a$. We apply the curve-tracing algorithm by solving the ODE (31)-(34), see Appendix D.3, for the function $H_{x,a}$ to find roots of the functions $H_{x,a}(\cdot,t)$ with $t\in[0,1]$. The root of $H_{x,a}(\cdot,1)$ gives $y=F(x)$.
@@ 4.2 Discontinuous semialgebraic functions
Discontinuities can arise in the ODE (7)-(8) since the value of $\dot{z}=M^{-1}b$ can be discontinuous across a boundary where $M$ is singular. Consider the simple example
@@
Every semialgebraic set has finitely many semialgebraic connected components, so every semialgebraic function is piecewise continuous on finitely many pieces. Our approach is to first show how we are able to represent the characteristic function of semialgebraic sets using the ODE (7)-(8). From these, we can represent a semialgebraic decomposition of the domain $\mathbb{R}^{m}$, and apply the continuous homotopy arguments above on each continuous region. We finally glue everything together in a way consistent with solving the ODE (7)-(8). The full proof is in Appendix E.
@@ 6 Discussion and future work
One interesting modification is to change line 3 of Algorithm 1 to have the ISD network $\mathcal{N}$ output an $LU$ factorization of $M$ rather than $M$ itself. This change would increase the speed of the networks since solving two triangular systems $LUz=b$ is significantly cheaper than solving a general linear system $Mz=b$. Furthermore, we could create a continuous variant of SANNs by requiring the diagonal elements of $L$ and $U$ to be positive and bounded away from 00. Such changes likely have many practical advantages; however, our expressivity proofs would no longer directly apply. Not every ISD matrix $M$ has an $LU$ decomposition in terms of ISD $L$ and $U$, and it is not clear whether this would be sufficient to apply our homotopy continuation arguments. Regardless, these modified SANNs compute a large class of semialgebraic functions not possible for most other architectures.
@@ Acknowledgments
The authors thank the anonymous ICLR reviewers for helpful comments on an ealier version of this manuscript. M. V. de Hoop and S. D. Mis acknowledge the support of the Simons Foundation under the MATH + X Program, the Department of Energy under grant DE-SC0020345, the Occidental Petroleum Corporation and the corporate members of the Geo-Mathematical Imaging Group at Rice University. M. L. was partially supported by a AdG project 101097198 of the European Research Council, Centre of Excellence of Research Council of Finland and the FAME flagship of the Research Council of Finland (grant 359186). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the funding agencies or the EU. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Lattice-ordered rings of piecewise polynomials
We use the following standard conventions: Group operations are denoted “$+$”, and ring multiplication operations are denoted ”*”. When there is no ambiguity, the ring multiplication operation $f*g$ may be written $fg$. If $R$ is a given ring, the polynomial ring in $n$ variables with coefficients in $R$ is denoted $R[X_{1},\dots,X_{n}]$. Direct products are denoted “$\times$”, and we reserve $\otimes$ for Kronecker products of matrices. Lattice operations are denoted $\wedge$ for ”inf” (”meet”) and $\vee$ for “sup” (“join”). We use the shorthand $f^{+}:=f\vee 0$, $f^{-}:=(-f)\vee 0$, and $|f|:=f\vee(-f)$. In this paper,
@@
In the terminology of ReLU networks, $f^{+}$ is precisely $\sigma(f)$ when the lattice is built over $\mathbb{R}^{n}$ and the meet/join operations are component-wise min/max.
@@
We will often refer to an $\ell$-ring $(R,+,*,\wedge,\vee)$ simply by it’s underlying set $R$.
@@
The original masters did not require $\ell$-rings to have commutative addition (i.e. be an Abelian groups) or have multiplicative identities. These generalizations are not important to us here, and what we call $\ell$-ring” is often called $\ell$-ring with unity” in the literature. All our $\ell$-rings are abelian groups and have multiplicative identity.
Every $\ell$-ring is a distributive lattice [^1] Only compatibility condition equation 17 is required, so in fact every so-called $\ell$-group is a distributive lattice. [Birkhoff, 1948]; that is,
@@
Garrett Birkhoff demonstrated $\ell$-rings can have surprising pathologies; for example, he constructed an $\ell$-ring whose multiplicative identity was both negative and a square [Birkhoff and Pierce, 1956]. The following definition introduces an additional compatibility condition that prevents the most troublesome pathologies.
@@
Condition 19 can be equivalently stated [Madden, 2011]
@@
  • Condition 17: Suppose $f\leq g$, which means $f(x)\leq g(x)$ for all $x\in D$:
@@
  • Condition 18: Suppose $f,g\geq 0$, meaning $f(x),g(x)\geq 0$ for all $x\in D$. Then $f(x)g(x)\geq 0$ since $R$ is an $\ell$-ring.
@@
  • Condition 19: Suppose $f,g,h\geq 0$ and $f\wedge h=0$, meaning $f(x),g(x),h(x)\geq 0$ and $f(x)\wedge h(x)=0$ for all $x\in D$. Then
@@
We need only verify the compatibility conditions equation 17, equation 18 and equation 19. But since these conditions hold in $R$, they clearly hold in $R_{1}$ as well. ∎
The requirement that $R_{1}\subseteq R$ form a sublattice in addition to a subring (rather than just a subring) is a small oversight in [Birkhoff and Pierce, 1956]. For example, polynomials form a subring of the $f$-ring of continuous real functions, but they are not a lattice ($x$ and $-x$ are both polynomials, but $x\vee-x=|x|$ is not), so polynomials do not themselves form an $f$-ring. Indeed, this fact will be relevant below to us below.
Definition 20 (Totally-ordered $\ell$ ring).
@@
def-totally-ordered-ell-ring An $\ell$-ring $R$ is *totally ordered* if both $f\wedge g$ and $f\vee g$ are in $\{f,g\}$ for all $f,g\in R$.
@@
  • Case $h=0$: From the $\ell$-ring compatibility condition equation 18, $fg\geq 0$ and $gf\geq 0$. Thus $fg\wedge h=fg\wedge 0=0$, likewise $gf\wedge h=gf\wedge 0=0.$
@@
This theorem first appeared in Henriksen and Isbell [1962]. The proof is surprisingly non-trivial, and while the key insights were provided by Henricksen and Isbell, they left the proof as an exercise to the reader. The first full proof appears to have been recorded in Hager and Johnson [2010].

### v2

@@ 1 Introduction
Many classical numerical algorithms compute semialgebraic functions—functions whose graphs are defined by finitely many polynomial equalities and inequalities. Such functions include anything computable using finitely many real variables and finitely many operations addition, subtraction, multiplication, division, extraction of roots, and branching if statements with polynomial decision boundaries. Familiar primitives from numerical linear algebra, such as solving linear systems, Schur complements, and extraction of real eigenvalues are also semialgebraic, as are the solution operators of optimization problems with polynomial objective functions and constraints (Press et al., 2007). Polynomial functions of $n$-th roots $n]{\cdot}$ are widely used in computation of eigenvalues of non-symmetric matrices and in the perturbation theory of eigenvalues of linear operators (Kato, 1995). Also, Finite Element approximations of partial differential equations lead to solving linear systems whose solutions are semialgebraic functions (Beltzer, 1990). Due to their ubiquity, it is likely that the space of semialgebraic functions provides good inductive bias for operator learning with neural networks. This ansatz leads to our central question: Is it possible to design a neural network architecture capable of exactly representing any semialgebraic function?
In this paper, we combine techniques from classical numerical analysis, namely polynomial homotopy continuation methods for root finding, with ReLU activation functions to create neural networks capable of exactly representing any bounded semialgebraic function. The starting point for our idea is the observation that, under mild conditions, the graph of a semialgebraic function $F$ given by the relation $F(x)=y$ can be encoded as the kernel of a continuous piecewise polynomial $G$, i.e., the set of $(x,y)$ where $G(x,y)=0$ (Section 2). Therefore, in principle, we can use a neural network combining polynomials and ReLU activations to learn $G$, then append a root-finding procedure such as Newton’s method to compute $F(x)=\texttt{root}(G(x,\cdot))=y$ in a manner similar to Bai et al. (2019). However, this straightforward approach has numerous practical difficulties: there is no guarantee such a neural network would have a root $y$ for every $x$, roots may not be unique, the root-finding procedure may never converge, etc. To remedy these issues, we take inspiration from homotopy continuation methods for root finding. We construct a function $H(x,y,s)$ that continuously deforms from a simple function $G_{0}$ to the target function $G$:
@@ 1.2 Our contribution
We introduce the SANN architecture and provide an exact characterization of the functions it represents. Our key contributions are as follows:
@@
We also argue SANNs are naturally suited to a wide variety of difficult problems, with applications to solving linear systems (Section 5), nonlinear inverse problems (Section F), general optimization problems (Section G.1), and transformers (Section G.3). We also demonstrate through numerical experiments that SANNs can be trained using standard techniques (Sections 5.2 and F.1), and discuss future research directions (Section 6).
@@ 2.1 Inf-sup definable piecewise polynomials
Our networks are constructed from the piecewise polynomials formed by combining polynomials with ReLU activations. Using these components, we can represent the pointwise minimum and maximum of any finite collection of polynomials. In mathematical literature, such structures are referred to as lattices. Instead of “min” and “max,” the terms “inf” and “sup” are traditionally used in lattice theory; since our constructions rely heavily on lattice theory, we adopt this terminology here. Notably, we do not use the word “lattice” to refer to a grid-like structure (e.g., $\mathbb{Z}^{d}$), but rather a set of functions that is closed under the “min” and “max” operations.
@@
def-isd Let $D\subseteq\mathbb{R}^{m}$. $ISD(D)$ is the lattice generated by the polynomials $\mathbb{R}[x_{1},\dots,x_{m}]$, viewed as functions $D\to\mathbb{R}$, together with the $\min$ and $\max$ operations. The vectors of $n$ $ISD(D)$ functions are denoted $ISD(D,\mathbb{R}^{n})$. For any non-negative integer $k$, $ISD^{k}(D,\mathbb{R}^{n})$ are the $k$-times differentiable functions in $ISD(D,\mathbb{R}^{n}).$
The shorthand “$f$ is $ISD$” means $f\in ISD(D,\mathbb{R}^{n})$ for some appropriate $D$ and $n$. Our nomenclature is taken from Mahé (1984) and Mahé (2007).
@@
It is currently unknown whether every continuous real piecewise polynomial is ISD; this is the famous Pierce–Birkhoff conjecture (Bochnak & Efroymson, 1980; Madden, 2011). ISD functions suffice for our purpose to build semialgebraic neural networks, regardless whether the Pierce–Birkhoff conjecture ultimately holds or not.
@@ 2.3 Semialgebraic functions as kernels of ISD functions
prop-semialgebraic-sets-as-kernels $S\subset\mathbb{R}^{m}$ is a closed semialgebraic set if and only if there exists $f\in ISD^{1}(\mathbb{R}^{m},\mathbb{R}_{\geq 0})$ such that $\ker(f)=S$, where $\ker(f):=\{x\in\mathbb{R}^{m}:f(x)=0\}$.
@@ 3.2 SANN architecture
[…]
@@
Matrix $M(x,z,s)$ and vector $b(x,z,s)$ are the output of an ISD network ${\mathcal{N}}(x,z,s)$. The function clamp-sol is defined
[…]
The function “clamp” operates component-wise on each element of a vector, and is defined
$\textrm{clamp}(a,low,high):=\begin{cases}a,&a\in[low,high]\\ low,&a<low\\ high&a>high.\end{cases}$ (5)
These functions guarantee the iteration (3) is well-defined even when $\mathcal{N}$ produces singular or nearly-singular output matrices $M$. Lastly, $\Pi$ projects onto the first $n$ components of its input.
The iteration (3) is the Euler finite-difference scheme that approximates the solution of the ODE
[…]
@@ 3.3 SANN architecture: Algorithmic presentation
In Algorithm 1, $\texttt{ODESolve}(\dot{z},z(0))$ outputs $z_{N}$ according to the finite difference scheme in formulas (2), or another numerical solver chosen by the programmer. This interface reflects popular libraries for numerically solving ODEs, e.g. Malengier et al. (2018); Kidger (2021).
@@
Our expressivity proofs below refer to the true solution $z(1)$ to the ODE (6)–(7) rather than the numerical solution $z_{N}.$ The numerical ODE solver may introduce approximation error into SANNs, even for semialgebraic functions, but this approximation error is of a fundamentally different character than that which normally arises in neural networks, and the mitigation strategies will be different for SANNs. For example, increasing the number of steps $N$ will decrease the approximation error without increasing the number of parameters in the network. On the other hand, increasing the number of parameters in $\mathcal{N}$ does not decrease the approximation error, as it generally would for other architectures. Furthermore, SANNs may be constructed using other numerical strategies besides the finite-difference scheme (2). Changing this scheme does not affect our expressivity proofs in this paper, but there may be other implications in practice.
@@ 3.4 Training
In the supervised learning regime, we are given training data $\{(x^{(\ell)},y^{(\ell)}):\ \ell=1,2,\dots,L\}$, $y^{(\ell)}=F(x^{(\ell)})$ where $F$ is the target semialgebraic function. The output of a SANN with input $x$ is $y_{N}(x):=\Pi z_{N}$ and the intermediate update directions given in equation (3) are $\dot{y}_{j}(x):=\Pi\dot{z}_{j}$. In our experiments, we found it helpful to include a “direction loss” term to train each $y_{j}$ for $j=1,\dots,N$:
@@ 4 Expressivity of SANNs
In this section, we show that SANNs are capable of exactly representing any bounded semialgebraic function. We start by showing that every continuous semialgebraic function can be evaluated by a particular homotopy continuation method that is computable by iteratively solving the ODE (6)–(7) with different parameter functions $\mathcal{N}$. We then extend this result to cover discontinuous semialgebraic functions by first separating each connected component of the graph, applying results for the continuous case to obtain homotopies that represent each piece, then glueing these homotopies together in a manner consistent with ODE (6)–(7) to obtain the SANN. Additional details, including background in semialgebraic geometry and full proofs of the theorems we present here, are in Appendices D and E.
@@ 4.1 Continuous semialgebraic functions
The next theorem shows the family of homotopies in equation (10) is capable of representing any continuous bounded semialgebraic function $F$ for almost every choice of $a$; more precisely every $a$ except possibly on a semialgebraic set of dimension less than $n$ (c.f. the Transversality theorem in D.1). The requirement that $F$ be bounded is always satisfied when the domain of continuous $F$ is compact, which is a common and reasonable assumption in the context of machine learning with finite training data. The proof of Theorem LABEL:thm-continuous-semialgebraic-functions-computable-by-homotopy is in appendix E.
@@
thm-continuous-semialgebraic-functions-computable-by-homotopy Let $F:\mathbb{R}^{m}\to U_{F}$ be a given continuous semialgebraic function. For every ${x\in\mathbb{R}^{m}}$ there exists $H_{x,a}$ with the form (10) such that for almost every ${a\in U_{F}}$, the following hold:
@@
Theorem LABEL:thm-continuous-semialgebraic-functions-computable-by-homotopy is used as follows: When $t=0$ the function $y\mapsto H_{x,a}(y,0)$ has the root $a$. We apply the curve-tracing algorithm by solving the ODE (30)–(33) (see Appendix D.3) for the function $H_{x,a}$ to find roots of the functions $H_{x,a}(\cdot,t)$ with $t\in[0,1]$. The root of $H_{x,a}(\cdot,1)$ gives $y=F(x)$.
@@ 4.2 Discontinuous semialgebraic functions
Discontinuities can arise in the ODE (6)–(7) since the value of $\dot{z}=M^{-1}b$ can be discontinuous across a boundary where $M$ is singular. Consider the simple example
@@
Every semialgebraic set has finitely many semialgebraic connected components, so every semialgebraic function is piecewise continuous on finitely many pieces. Our approach is to first show how we are able to represent the characteristic function of semialgebraic sets using the ODE (6)–(7). From these, we can represent a semialgebraic decomposition of the domain $\mathbb{R}^{m}$, and apply the continuous homotopy arguments above on each continuous region. We finally glue everything together in a way consistent with solving the ODE (6)–(7). The full proof is in Appendix E.
@@ 6 Discussion and future work
One interesting modification is to change line 3 of Algorithm 1 to have the ISD network $\mathcal{N}$ output an $LU$ factorization of $M$ rather than $M$ itself. This change would increase the speed of the networks since solving two triangular systems $LUz=b$ is significantly faster than solving a general linear system $Mz=b$. Furthermore, we could create a continuous variant of SANNs by requiring the diagonal elements of $L$ and $U$ to be positive and bounded away from 00. Such changes likely have many practical advantages; however, our expressivity proofs would no longer directly apply. Not every ISD matrix $M$ has an $LU$ decomposition in terms of ISD $L$ and $U$, and it is not clear whether this would be sufficient to apply our homotopy continuation arguments. Regardless, these modified SANNs compute a large class of semialgebraic functions not possible for most other architectures.
@@ Acknowledgments
The authors thank the anonymous ICLR reviewers for helpful comments on an ealier version of this manuscript. M. V. de Hoop and S. D. Mis acknowledge the support of the Simons Foundation under the MATH + X Program, the Department of Energy under grant DE-SC0020345, the Occidental Petroleum Corporation and the corporate members of the Geo-Mathematical Imaging Group at Rice University. M. Lassas was partially supported by a AdG project 101097198 of the European Research Council, Centre of Excellence of Research Council of Finland and the FAME flagship of the Research Council of Finland (grant 359186). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the funding agencies or the EU. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Lattice-ordered rings of piecewise polynomials
We use the following standard conventions: Group operations are denoted “$+$”, and ring multiplication operations are denoted “*”. When there is no ambiguity, the ring multiplication operation $f*g$ may be written $fg$. If $R$ is a given ring, the polynomial ring in $n$ variables with coefficients in $R$ is denoted $R[X_{1},\dots,X_{n}]$. Direct products are denoted “$\times$”, and we reserve $\otimes$ for Kronecker products of matrices. Lattice operations are denoted $\wedge$ for “inf” (“meet”) and $\vee$ for “sup” (“join”). We use the shorthand $f^{+}:=f\vee 0$, $f^{-}:=(-f)\vee 0$, and $|f|:=f\vee(-f)$. In this paper,
@@
In the terminology of neural networks, $f^{+}$ is precisely $\mathrm{ReLU}(f)$ when the lattice is built over $\mathbb{R}^{n}$ and the meet/join operations are component-wise min/max.
@@
We will often refer to an $\ell$-ring $(R,+,*,\wedge,\vee)$ simply by its underlying set $R$.
@@
The original masters did not require $\ell$-rings to have commutative addition (i.e. be an Abelian groups) or have multiplicative identities. These generalizations are not important to us here, and what we call “$\ell$-ring” is often called “$\ell$-ring with unity” in the literature. All our $\ell$-rings are abelian groups and have multiplicative identity.
Every $\ell$-ring is a distributive lattice [^1] Only compatibility condition (16) is required, so in fact every so-called $\ell$-group is a distributive lattice. (Birkhoff, 1948); that is,
@@
Garrett Birkhoff demonstrated $\ell$-rings can have surprising pathologies; for example, he constructed an $\ell$-ring whose multiplicative identity was both negative and a square (Birkhoff & Pierce, 1956). The following definition introduces an additional compatibility condition that prevents the most troublesome pathologies.
@@
Condition (18) can be equivalently stated (Madden, 2011)
@@
  • Condition (16): Suppose $f\leq g$, which means $f(x)\leq g(x)$ for all $x\in D$:
@@
  • Condition (17): Suppose $f,g\geq 0$, meaning $f(x),g(x)\geq 0$ for all $x\in D$. Then $f(x)g(x)\geq 0$ since $R$ is an $\ell$-ring.
@@
  • Condition (18): Suppose $f,g,h\geq 0$ and $f\wedge h=0$, meaning $f(x),g(x),h(x)\geq 0$ and $f(x)\wedge h(x)=0$ for all $x\in D$. Then
@@
We need only verify the compatibility conditions (16), (17) and (18). But since these conditions hold in $R$, they clearly hold in $R_{1}$ as well. ∎
The requirement that $R_{1}\subseteq R$ form a sublattice in addition to a subring (rather than just a subring) is a small oversight in (Birkhoff & Pierce, 1956). For example, polynomials form a subring of the $f$-ring of continuous real functions, but they are not a lattice ($x$ and $-x$ are both polynomials, but $x\vee-x=|x|$ is not), so polynomials do not themselves form an $f$-ring. Indeed, this fact will be relevant below to us below.
Definition 20 (Totally ordered $\ell$ ring).
@@
def-totally-ordered-ell-ring An $\ell$-ring $R$ is totally ordered if both $f\wedge g$ and $f\vee g$ are in $\{f,g\}$ for all $f,g\in R$.
@@
  • Case $h=0$: From the $\ell$-ring compatibility condition (17), $fg\geq 0$ and $gf\geq 0$. Thus $fg\wedge h=fg\wedge 0=0$, likewise $gf\wedge h=gf\wedge 0=0.$
@@
This theorem first appeared in Henriksen & Isbell (1962). The proof is surprisingly non-trivial, and while the key insights were provided by Henricksen and Isbell, they left the proof as an exercise to the reader. The first full proof appears to have been recorded in Hager & Johnson (2010).